simplify the expression
5^3/2*5^1/2

Answer by stanbon(75887)

5^3/2*5^1/2
--------------
= 5^((3/2)+(1/2))
---------
= 5^(4/2)
---
= 5^2
----
= 25
===========
Cheers,
Stan H.